Determine whether the lines are parallel, perpendicular or neither. There is a discrepancy three of us are having! lol)
3x + 6y =21
-6x + 3y = 9

Answer by Alan3354(69443)

Determine whether the lines are parallel, perpendicular or neither. There is a discrepancy three of us are having! lol)
3x + 6y =21
-6x + 3y = 9
------------
Solve for y
3x + 6y =21
y = (-1/2)x + 7/2
Slope m1 = -1/2
-------------------
-6x + 3y = 9
y = 2x + 3
Slope m2 = 2
-----
m1*m2 = -1 --> perpendicular
--------
If the slopes are negative inverses --> perpendicular
If the slopes are equal --> parallel